The most recent financial statements for Bello Co. are shown here: Income Statement Balance Sheet Sales $ 18,900 Current assets $ 11,700 Debt $ 15,700 Costs 12,800 Fixed assets 26,500 Equity 22,500 Taxable income $ 6,100 Total $ 38,200 Total $ 38,200 Taxes (21%) 1,281 Net income $ 4,819 Assets and costs are proportional to sales. Debt and equity are not. The company maintains a constant 30 percent dividend payout ratio. What is the internal growth rate? (Do not round intermediate calculations and enter your answer as a percent rounded 2 decimal places, e.g., 32.16.)
Answer #1

ROA=net income/total assets

=(4819/38200)=0.126151832

Retention ratio=1-dividend payout ratio

=(1-0.3)=0.7

Internal growth rate=(ROA*Retention ratio)/[1-(ROA*Retention ratio)]

(0.126151832*0.7)/[1-(0.126151832*0.7)]

which is equal to

=9.69%(Approx).